# Cross-Domain Underwater Image Enhancement Guided by No-Reference Image Quality Assessment: A Transfer Learning Approach

### v1

@@ Abstract
Single underwater image enhancement (UIE) is a challenging ill-posed problem, but its development is hindered by two major issues: (1) The labels in underwater reference datasets are pseudo labels, relying on these pseudo ground truths in supervised learning leads to domain discrepancy. (2) Underwater reference datasets are scarce, making training on such small datasets prone to overfitting and distribution shift. To address these challenges, we propose Trans-UIE, a transfer learning-based UIE model that captures the fundamental paradigms of UIE through pretraining and utilizes a dataset composed of both reference and non-reference datasets for fine-tuning. However, fine-tuning the model using only reconstruction loss may introduce confirmation bias. To mitigate this, our method leverages no-reference image quality assessment (NR-IQA) metrics from above-water scenes to guide the transfer learning process across domains while generating enhanced images with the style of the above-water image domain. Additionally, to reduce the risk of overfitting during the pretraining stage, we introduce Pearson correlation loss. Experimental results on both full-reference and no-reference underwater benchmark datasets demonstrate that Trans-UIE significantly outperforms state-of-the-art methods.
@@ 2.1 Underwater Image Restoration Methods
With the advancement of deep learning, data-driven underwater image enhancement methods have emerged and demonstrated strong competitiveness. In the early stages, due to the lack of paired training samples, researchers introduced generative adversarial networks (GANs) to improve underwater image quality, such as WaterGAN [31], FUnIE [21], UGAN [10], and UIE-DAL [48]. The release of underwater image datasets such as UIEB [29] and LSUI [40] has facilitated the development of more complex network architectures to optimize image enhancement performance, including WaterNet [29], Ucolor [30], TSDA [23], and Ushape [40]. These models incorporate various innovative techniques to better adapt to complex underwater scenarios. For example, Li et al. [30] proposed a transmission map-guided network that combines medium transmission maps with multi-channel inputs to enhance model robustness, while PUIE-Net [12] integrates a conditional variational autoencoder with adaptive instance normalization for adjustable image enhancement. Additionally, Ushape [40] and RCTNet [25] introduce Transformer-based architectures and attention mechanisms, significantly improving enhancement performance under complex underwater lighting and color variations.
@@ 3.1 Transfer Learning Framework
Our transfer learning framework is illustrated in Fig. 2. First, a pre-trained model, UIR-Net, is trained on a reference dataset by minimizing the pixel-level loss $\mathcal{L}_{\text{pix}}$ and the Pearson correlation loss $\mathcal{L}_{\text{cor}}$ to learn the mapping from degraded images to pseudo labels $I^{*}$, thereby obtaining a basic enhancement network. Next, UIR-Net is used to generate pseudo labels for the fine-tuning dataset, providing an initial mapping for the fine-tuning stage. For quality assessment of non-reference images, [20] indicates that the deep learning-based MUSIQ [24] is highly reliable; thus, we use MUSIQ [24] as our above-water NR-IQA to score the generated pseudo labels, obtaining the initial reference quality score $Q_{\text{reference}}$.
@@
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ denote the weight coefficients for the pixel-level loss, VGG [47] perceptual loss, and NR-IQA score loss. $Q(\hat{I})$ is the score assigned to the enhanced image $\hat{I}$ by NR-IQA within the wrapper. $Q_{\text{reference}}$ is the reference quality score obtained by evaluating the pseudo labels generated by UIR-Net using NR-IQA, it is responsible for adjusting the learning rate of $L_{\text{MUSIQ}}$ and does not participate in gradient computation.
@@ 3.2 Domain Discrepancy and Distribution Shift
We design the overall loss function as Eq. 1. If only the losses $L_{\text{pix}}$ and $L_{\text{vgg}}$ are used, the model may learn features that are biased by $\epsilon_{\text{pseudo}}$, causing the output feature $\varphi(\hat{I})$ to deviate from the true image feature $\varphi(I_{\text{real}})$. The introduction of the NR-IQA score loss $L_{\text{MUSIQ}}$ helps to align the quality of the output images with the target $\varphi(I_{\text{real}})$, thereby partially mitigating the domain discrepancy induced by $\epsilon_{\text{pseudo}}$.
During training, $L_{\text{total}}$ gradually converges, while $Q(\hat{I})$ increases. We assume that the mapping function $f(\cdot)$ maps the deep features of the image to the NR-IQA score space. When the model output $\hat{I}$’s feature approaches the above-water image feature $\varphi(I_{\text{real}})$, $Q(\hat{I})$ will tend to reach its maximum value. Hence, the total loss can be approximated by:
@@
During fine-tuning, we employ a training set that mixes reference and non-reference images and generate pseudo labels using a pre-trained model. The pixel-level loss $L_{\text{pix}}$ and the perceptual loss $L_{\text{vgg}}$ encourage the expected features of the generated image $\hat{I}$ to approach $\mu_{R}$. Simultaneously, the NR-IQA score loss $L_{\text{MUSIQ}}$ further enhances the quality score $E[Q(\hat{I})]$, driving the model to suppress noise during generation. In this way, the discrepancy between the reference noise $\epsilon_{R}$ and the non-reference noise $\epsilon_{N}$ is effectively reduced, thereby compensating for the feature distribution shift $\Delta_{\text{feat}}$.
@@
$L_{\text{total}}\approx\lambda\|\varphi(\hat{I})-\mu_{R}\|_{2}^{2}+\lambda_{3}\left(E[Q(\hat{I})]_{\text{desired}}-E[Q(\hat{I})]\right),$ (13)
@@ 3.3 Pretrained Network UIR-Net
In this study, we propose UIR-Net, a U-shaped Transformer architecture that captures global context and fine-grained details, providing a solid foundation for transfer learning. The framework is shown in Fig. 3. The input image $I\in\mathbb{R}^{H\times W\times C}$ is divided into overlapping patches using a $3\times 3$ convolution. The encoder reduces the spatial resolution through an Unshuffle operation while extracting global features with Transformer modules. The decoder restores the resolution using a Shuffle operation. Skip connections fuse low-level encoder features with high-level decoder semantics.
@@ 3.4 Pearson Correlation Loss
where $\hat{I}(i)$ denotes the $i$-th pixel of the image, and $\mu(\hat{I})$ and $\sigma(\hat{I})$ represent the mean and standard deviation of $\hat{I}$ (similarly for $I_{gt}$). The value of $\rho(\hat{I},I_{gt})$ lies in the range $[-1,1]$; when the two images are perfectly positively correlated, $\rho(\hat{I},I_{gt})=1$, and when they are negatively correlated, $\rho(\hat{I},I_{gt})=-1$.
@@ 4.4 Ablation Studies
Analysis of Transfer Learning Effectiveness. We compared four methods: (a) base: the pre-trained model without fine-tuning; (b) Trans-base: fine-tuning with only the pixel loss $L_{\text{pix}}$; (c) Trans-base+$L_{\text{vgg}}$: fine-tuning with pixel loss $L_{\text{pix}}$ and perceptual loss $L_{\text{vgg}}$; and (d) Trans-UIE: fine-tuning with all losses.
Comparisons were conducted on both reference and non-reference datasets, with quantitative results presented in Tab. 3. Our complete solution consistently outperforms the others. Comparing base and Trans-base, we observe that Trans-base, which uses pseudo-labels as training supervision, may amplify existing biases or errors, leading to confirmation bias. The introduction of perceptual loss $L_{\text{vgg}}$ in Trans-base+$L_{\text{vgg}}$ effectively mitigates this bias. Upon observing Trans-UIE, we note a decrease in PSNR scores on LSUI. This can be attributed to the introduction of evaluation loss $L_{\text{MUSIQ}}$, which reduce the domain gap between underwater and real surface images. Qualitative results, shown in Fig. 6, highlight that Trans-UIE delivers the best image restoration, with superior color fidelity and rich detail.

### v2

@@ Abstract
Single underwater image enhancement (UIE) is a challenging ill-posed problem, and its development is hindered by two main issues: (1) The labels in underwater reference datasets are pseudo ground truths with severe color distortions; relying on them in supervised learning leads to domain discrepancy. (2) Underwater reference datasets are scarce, and training on such small datasets tends to cause overfitting and distribution shifts from real underwater conditions. To address these challenges, we propose Trans-UIE, a transfer learning-based UIE method. Our approach captures the fundamental paradigms of UIE through pretraining and utilizes a hybrid dataset for fine-tuning, thereby reducing the distribution gap from real underwater environments. Moreover, our method incorporates no-reference image quality assessment (NR-IQA) metrics from above-water scenes to guide the cross-domain transfer learning process, mitigating the color distortions introduced by pseudo labels and narrowing the domain gap between underwater and above-water images, while avoiding confirmation bias. Additionally, to reduce the risk of overfitting during the pretraining phase, we introduce a Pearson correlation loss. Experimental results on both full-reference and no-reference underwater image benchmark datasets demonstrate that Trans-UIE achieves outstanding performance in image enhancement quality.
@@ 2.1 Underwater Image Restoration Methods
With the advancement of deep learning, data-driven underwater image enhancement methods have emerged and demonstrated strong competitiveness. In the early stages, due to the lack of paired training samples, researchers introduced generative adversarial networks (GANs) to improve underwater image quality, such as WaterGAN [32], FUnIE [22], UGAN [10], and UIE-DAL [49]. The release of underwater image datasets such as UIEB [30] and LSUI [41] has facilitated the development of more complex network architectures to optimize image enhancement performance, including WaterNet [30], Ucolor [31], TSDA [24], and Ushape [41]. These models incorporate various innovative techniques to better adapt to complex underwater scenarios. For example, Li et al. [31] proposed a transmission map-guided network that combines medium transmission maps with multi-channel inputs to enhance model robustness, while PUIE-Net [12] integrates a conditional variational autoencoder with adaptive instance normalization for adjustable image enhancement. Additionally, Ushape [41] and RCTNet [26] introduce Transformer-based architectures and attention mechanisms, significantly improving enhancement performance under complex underwater lighting and color variations. The most closely related work to this paper is Semi-UIR [21], which innovatively adopts a semi-supervised learning approach. By leveraging MUSIQ [25] scores for pseudo-label selection, it incorporates non-reference datasets into training and achieves impressive results. However, this method primarily focuses on mitigating overfitting on small reference datasets and does not explicitly address or analyze the bias between pseudo labels and above-water images. Moreover, the performance of its single-stage training strategy is prone to limitations caused by confirmation bias.
@@ 3.1 Transfer Learning Framework
Our transfer learning framework is illustrated in Fig. 2. First, a pre-trained model, UIR-Net, is trained on a reference dataset by minimizing the pixel-level loss $\mathcal{L}_{\text{pix}}$ and the Pearson correlation loss $\mathcal{L}_{\text{cor}}$ to learn the mapping from degraded images to pseudo labels $I^{*}$, thereby obtaining a basic enhancement network. Next, UIR-Net is used to generate pseudo labels for the fine-tuning dataset, providing an initial mapping for the fine-tuning stage. we use MUSIQ [25] as our above-water NR-IQA to score the generated pseudo labels, obtaining the initial reference quality score $Q_{\text{reference}}$.
@@
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ denote the weight coefficients for the pixel-level loss, VGG [48] perceptual loss, and NR-IQA score loss. Q(I^)Q($\hat${I}) is the score assigned to the enhanced image $\hat{I}$ by NR-IQA within the wrapper. $Q_{\text{reference}}$ is the reference quality score obtained by evaluating the pseudo labels generated by UIR-Net using NR-IQA, it is responsible for adjusting the learning rate of $L_{\text{MUSIQ}}$ and does not participate in gradient computation.
3.2 Reliable Metric Selection
We expect a suitable no-reference image quality assessment (NR-IQA) metric to meet two key criteria: (1) it should accurately reflect the quality changes of underwater images before and after enhancement, and (2) it should exhibit strong evaluation capability for above-water images. To identify the most appropriate NR-IQA metric, we conduct an empirical analysis of several existing NR-IQA methods.
For the first criterion, we start with a set of degraded images and a set of clean images. By linearly mixing the two sets at different ratios, we generate a series of images with varying quality levels. As the proportion of degraded content increases, the visual quality of the mixed images gradually declines. Therefore, a reliable NR-IQA metric is expected to produce scores that monotonically decrease with image degradation. Metrics that satisfy this monotonicity law are considered to have strong assessment capability. For the second criterion, we evaluate each method by comparing its PLCC score on a standard above-water dataset.
Following these principles, we evaluate six NR-IQA methods on the EUVP benchmark dataset [23], which covers a wide range of underwater scenes, and the above-water dataset KonIQ [20]. As shown in Fig. 3, the deep learning-based MUSIQ [25] shows the best consistency with the monotonicity law and is therefore selected to assess the reliability of network outputs.
Figure 3: The results of different non-reference IQA indicators on EUVP [23] and KonIQ [20], including UIQM [39], UCIQE [51], BRISQUE [20], NIQE [37], PaQ2PiQ [20] and MUSIQ [25].
@@ 3.3 Domain Discrepancy and Distribution Shift
[…]
During training, $L_{\text{total}}$ gradually converges, while Q(I^)Q($\hat${I}) increases. We assume that the mapping function $f(\cdot)$ maps the deep features of the image to the NR-IQA score space. When the model output $\hat{I}$’s feature approaches the above-water image feature $\varphi(I_{\text{real}})$, Q(I^)Q($\hat${I}) will tend to reach its maximum value. Hence, the total loss can be approximated by:
@@
During fine-tuning, we employ a training set that mixes reference and non-reference images and generate pseudo labels using a pre-trained model. The pixel-level loss $L_{\text{pix}}$ and the perceptual loss $L_{\text{vgg}}$ encourage the expected features of the generated image $\hat{I}$ to approach $\mu_{R}$. Simultaneously, the NR-IQA score loss $L_{\text{MUSIQ}}$ further enhances the quality score E[Q(I^)]E[Q($\hat${I})], driving the model to suppress noise during generation. In this way, the discrepancy between the reference noise $\epsilon_{R}$ and the non-reference noise $\epsilon_{N}$ is effectively reduced, thereby compensating for the feature distribution shift $\Delta_{\text{feat}}$.
@@
[…]
@@ 3.4 Pretrained Network UIR-Net
In this study, we propose UIR-Net, a newly designed architecture built upon the Restormer [56] backbone, which effectively captures both global context and fine-grained details, providing a solid foundation for transfer learning. The framework is shown in Fig. 4. The input image $I\in\mathbb{R}^{H\times W\times C}$ is divided into overlapping patches using a $3\times 3$ convolution. The encoder reduces the spatial resolution through an Unshuffle operation while extracting global features with Transformer modules. The decoder restores the resolution using a Shuffle operation. We introduce skip connections to fuse low-level encoder features with high-level decoder semantics.
@@ 3.5 Pearson Correlation Loss
[…]
@@ 4.4 Ablation Studies
[…]